OpenStudy (anonymous):

Which point is the vertex of the parabola y = x2+4x+1? A. (-2, -3) B. (2, 3) C. (-2, 3) D. (2, -3)

OpenStudy (anonymous):

Agreed^

OpenStudy (anonymous):

thanks :P

hero (hero):

A only

OpenStudy (anonymous):

??

OpenStudy (anonymous):

how many vertices does this one poor quadratic have?

hero (hero):

satellite just post your solution

OpenStudy (anonymous):

dunno. pick all 4 because apparently there can be more than one

OpenStudy (radar):

Nah,vertex is singular, plural would be vertices lol

hero (hero):

Once again, that's not a solution sat

hero (hero):

I'm sticking with A = (-2,-3)

OpenStudy (anonymous):

I agree with Hero (A)

OpenStudy (anonymous):

@belaski \[x=-\frac{b}{2a}=-\frac{4}{2\times 1}=-2\]is the first coordinate second coordinate is \[(-2)^2+4(-2)+1=4-8+1=-3\]

OpenStudy (anonymous):

well your only suppose to chose 1

hero (hero):

satellite just didn't want to give me any credit

OpenStudy (anonymous):

so answer is \[(-2,-3)\] trust me, i am a carnival barker

OpenStudy (anonymous):

ok well thanks guys

hero (hero):

Mancake, prove that your original answer is correct. Only then will I believe you to be a professor.

OpenStudy (anonymous):

I think prof. mancake is right

OpenStudy (anonymous):

I just used the same formulas and i got the same answers

OpenStudy (anonymous):

A. (-2, -3) y = x2+4x+1 h = -4/2 = -2 k = (-2)^2 -4* 2 + 1 = -3